Pipe A can fill a tank 1.25 times faster than pipe B. When both pipes are opened, they
fill the tank in five hours. How long would it take to fill the tank if only pipe B is used?Pipe A can fill a tank 1.25 times faster than pipe B. When both pipes are opened, they
fill the tank in five hours. How long would it take to fill the tank if only pipe B is used?Pipe A can fill a tank 1.25 times faster than pipe B. When both pipes are opened, they
fill the tank in five hours. How long would it take to fill the tank if only pipe B is used?

Answer by nerdybill(6962)

Pipe A can fill a tank 1.25 times faster than pipe B. When both pipes are opened, they fill the tank in five hours. How long would it take to fill the tank if only pipe B is used?
Let x = time (hours) it takes for just pipe B
then
1.25x = time (hours) it takes for just pipe A
.
5(1/x + 1/(1.25x)) = 1
multiplying both sides by 1.25x:
5(1.25 + 1) = 1.25x
5(2.25) = 1.25x
11.25 = 1.25x
11.25/1.25 = x
9 hours = x